OpenStudy (anonymous):

A car journey of 380 km lasts 4 hours. Part of this is on a motorway at an average speed of 110kmh^-1, the rest on country roads at an average speed of 70kmh^-1. (i) Write this information as pair of simultaneous equations. (ii) Solve you equations to find how many kilometres of the journey is spent on each type of road. So, I know what the answers are but I can't figure out how to get to it myself. Please will you help by showing me how?

OpenStudy (welshfella):

use the relation v = d/t v= speed , d = distance travel on a motorway and t = time taken on the motorway so for motorway travel we have 110 = d/t and for the country roads 70 = (380 - d) / (4 - t)

OpenStudy (welshfella):

so there are your 2 equations

OpenStudy (welshfella):

writing the equation without fractions we get d = 110t 380 - d = 70(4 - t) to solve for t substitute d = 110t in the second equation 380 - 110t = 280 - 70t 380 - 280 = 110t - 70t can you continue with this?

OpenStudy (welshfella):

oh they want number of kilometers on each type of road I should have solved for d no matter when we find t we can then plug in the values to find the distances

OpenStudy (welshfella):

fo you follow it OK so far?

OpenStudy (welshfella):

are you there?

OpenStudy (anonymous):

Yes, I am just a bit confused since those are not the answers my book got? I will give those too you now...

OpenStudy (anonymous):

t(subscript 1) + t(subscript 2) =4; 110t(subscript 1) + 70t(subscript 2) = 380 Sorry I cant type the subscript numbers on my computer.

OpenStudy (anonymous):

That is what my textbook has said are the answers to the first question.

OpenStudy (welshfella):

yes thats basically the same as i have bu they have called the two times t1 and t2 whereas ive called them t and 4 - t because total time was 4 hours

OpenStudy (welshfella):

if you solve for t you;ll find t = time for motoirway journey = 2.5 hours and so time foe road journey = 4-2.5 = 1.5 and distance for motorway = 110 * 2.5 = 275km and for raoad = 70 8 1.5 = 105 km

OpenStudy (welshfella):

t1 + t1 = 4 110 t1 + 70 t2 = 380 is what your book has said

OpenStudy (welshfella):

do you want me to show you how to solve this?

OpenStudy (welshfella):

there is often a few different ways to solve problems like this.

OpenStudy (anonymous):

Oh ok. I think I understand now- thank you very much!

OpenStudy (welshfella):

oh that should be t1 + t2 = 4

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

t1 being time for motorway and t2 time for roads

OpenStudy (anonymous):

Ok. You have helped me a lot!

OpenStudy (welshfella):

yw